OpenStudy (anonymous):

4x^4=-29x^2-25=0 solve the equation with rational exponents

OpenStudy (saifoo.khan):

You equation contain 2 equal signs

OpenStudy (anonymous):

sorry disregard the =0

OpenStudy (saifoo.khan):

Start with: x^2 = y Rewrite: 4y^2 = -29y -25 Solve the quadratic.

OpenStudy (anonymous):

y=-1, -25/4 ?

OpenStudy (saifoo.khan):

Yes. So now we have some imaginary roots.

OpenStudy (saifoo.khan):

Insert the values back in: x^2 = y